OpenStudy (anonymous):

solve by completing the square.. x^2+14x=10

OpenStudy (anonymous):

x^2 + 14x + 49 = 10 + 49 (x+7)^2 = 59 x + 7 = pm sqrt(59) x = -7 pm sqrt(59)

OpenStudy (anonymous):

thanks;)